OpenStudy (jewelliasmith1):

an electronics manufacturer recently created a new version of a popular device. it also created this function to represent the profit, P(x), in tens of thousands of dollars, that the company will earn based on manufacturing x thousand devices: P(x)=-0.16x^2+21.6x-400. The profit functions for the first version of the device was similar to the profit function for the new version. As a matter of fact, the profit function for the first version is a transformation of the profit function for the new version. for the value x=40, the original profit function is half the size of the new profit functi-

OpenStudy (jewelliasmith1):

-ion. write two function transformations in terms of P(x) that could represent the original profit function?

OpenStudy (sachintha):

@mathmate

OpenStudy (mathmate):

@jewelliasmith1 Have you finished reading the question?

OpenStudy (jewelliasmith1):

yes @mathmate

OpenStudy (mathmate):

Good! This problem is expressed in real-life application form, which makes it look like it is long. The mathematical part can be resumed into much simpler terms. Would you like to explain to me the mathematical aspect of the question?

OpenStudy (jewelliasmith1):

im so confused

OpenStudy (mathmate):

Can you at least extract what is given to you from the question? Whatever comes to your mind? This practice will help you solve other word problems later on.

OpenStudy (jewelliasmith1):

p(40)=-0.16(40)^2+1.6(40)-400

OpenStudy (mathmate):

That's good! What does this function represent?

OpenStudy (jewelliasmith1):

the device thingy

OpenStudy (mathmate):

By the way, the function is p(40)=-0.16(40)^2+\(\color{red}{2}\)1.6(40)-400

OpenStudy (mathmate):

It is the amount of profit when 40 devices are sold. Reread the question and check if what I said is true. I make mistakes too! lol

OpenStudy (jewelliasmith1):

yea

OpenStudy (mathmate):

Good! Now, recall that P(x) represents the profit for selling x units with the \(new\) device. The old device has a profit function (I will call it PT(x)) which is a transformation of P(x), right?

OpenStudy (jewelliasmith1):

i guess

OpenStudy (mathmate):

Reread the question to make sure you understand the same way as I said earlier. Otherwise you'd be following me blindly, and you won't be able to solve your next problem!

OpenStudy (jewelliasmith1):

ok

OpenStudy (jewelliasmith1):

so what do i put as the answer

OpenStudy (mathmate):

You'll need to do more before you get the answer! If that's understood, so far, you need to decide on a transformation of P(x) to PT(x). Two transformations are possible, 1. you can do a horizontal stretch, i.e. PT(x)=P(kx) with the stretch factor k, or 2. you can do a horizontal translation, i.e. PT(x)=P(x-b) where k and b are constants to be determined. We don't know what they are, yet.

OpenStudy (jewelliasmith1):

im so confused

OpenStudy (mathmate):

That is why you need to understand the question as a start. Please reread the question, especially the part: The profit function for the first version of the device was similar to the profit function for the new version. As a matter of fact, the profit function for the first version is a transformation of the profit function for the new version.

OpenStudy (mathmate):

Is it a little less confusing now? The question explains that the profit function for the old product is a transformation of that of the new one, hence PT(x)=transformation of P(x) And we can try the transformation vertical stretch, or horizontal translation, both are possible.

OpenStudy (mathmate):

Perhaps reading the following will help to clarify a little: http://www.purplemath.com/modules/fcntrans.htm

OpenStudy (jewelliasmith1):

oh ur talking about graphs

OpenStudy (mathmate):

Yes, that too! Actually, the graph of P(x) is as follows: http://prntscr.com/bg9ow2

OpenStudy (jewelliasmith1):

ok

OpenStudy (mathmate):

To proceed, we will now make use of the information: For the value x=40, the original profit function is half the size of the new profit functiion. which can be expressed in a mathematical equation: PT(40)=(1/2) P(40). Does that make sense to you?

OpenStudy (jewelliasmith1):

yea

OpenStudy (mathmate):

Good, so let's start with the vertical stretch as a transformation. A vertical stretch is simply PT(x)=a P(x), where a is a constant. Putting in PT(40)=aP(40)=(1/2)P(40), can you solve for a? hint: use the second part: aP(40)=(1/2)P(40) to solve for a.

OpenStudy (jewelliasmith1):

i dont get it

OpenStudy (mathmate):

The question tells us: "As a matter of fact, the profit function for the first version is a transformation of the profit function for the new version. " So we're making a transformation (of any kind, but we chose the simplest kind, a vertical stretch) to satisfy the other requirement: " For the value x=40, the original profit function is half the size of the new profit functiion." Therefore we come up with the previous equation: PT(40)=aP(40)=(1/2)P(40), can you solve for a? hint: use the second part: aP(40)=(1/2)P(40) to solve for a.

OpenStudy (jewelliasmith1):

ok

OpenStudy (mathmate):

So now you need to solve for "a" in the equation in order to complete the definition of the transformation.

OpenStudy (jewelliasmith1):

ok

OpenStudy (jewelliasmith1):

5.2

OpenStudy (mathmate):

Not quite. The equation to solve is \(\large a~ P(40)=(1/2)~P(40)\) we need the value of "a". Try to transpose or cancel common factors.

OpenStudy (mathmate):

Perhaps this might help you! https://www.khanacademy.org/math/k-8-grades/cc-eighth-grade-math/cc-8th-solving-equations